# mCLM: A Modular Chemical Language Model that Generates Functional and Makeable Molecules

### v2

@@ Abstract
Despite their ability to understand chemical knowledge, large language models (LLMs) remain limited in their capacity to propose novel molecules with desired functions (e.g., drug-like properties). In addition, the molecules that LLMs propose can often be challenging to make, and are almost never compatible with automated synthesis approaches. To better enable the discovery of functional small molecules, LLMs need to learn a new molecular language that is more effective in predicting properties and inherently synced with automated synthesis technology. Current molecule LLMs are limited by representing molecules based on atoms. In this paper, we argue that just like tokenizing texts into meaning-bearing (sub-)word tokens instead of characters, molecules should be tokenized at the level of functional building blocks, i.e., parts of molecules that bring unique functions and serve as effective building blocks for real-world automated laboratory synthesis. This motivates us to propose mCLM, a modular Chemical-Language Model that comprises a bilingual language model that understands both natural language descriptions of functions and molecular blocks. mCLM front-loads synthesizability considerations while improving the predicted functions of molecules in a principled manner. Experiments on 430 FDA-approved drugs showed that mCLM is capable of significantly improving chemical functions critical to determining drug potentials. mCLM, with only 3B parameters, also achieves improvements in synthetic accessibility relative to 7 other leading generative AI methods including GPT-5. When tested on 122 out-of-distribution medicines using only building blocks/tokens that are compatible with automated modular synthesis, mCLM outperforms all baselines in property scores and synthetic accessibility. mCLM can also reason on multiple functions and iteratively self-improve to rescue drug candidates that failed late in clinical trials (“fallen angels”).  [^1] All codes, data, and models will be released publicly upon publication.
@@ 1 Introduction
Therefore, our goal is not to make all drug-like molecules, rather, our goal is to figure out how to make the right ones, better, faster, stronger. Unlike machines, human scientists are inherently “multilingual,” seamlessly navigating diverse modalities—from natural language and scientific figures in literature to complex scientific data such as molecular structures in knowledge bases. In contrast, most prior work on molecule discovery trains large language models on a single modality. Moreover, Human scientists “think before they talk,” grounding their reasoning in deeply reflective and deliberate reflection and critical evaluation to generate new hypotheses. Current models lack this critical thinking capacity, limiting their ability to contribute meaningfully to discovery. In particular, the human body is a highly complex, interconnected system, and drug discovery is essentially a multi-objective optimization problem. It involves balancing factors such as drug absorption, first-pass metabolism, bioavailability, distribution, protein binding, and clearance. However, improving one property often comes at the expense of another. Many promising drug candidates have failed in the final stages of FDA approval due to an unacceptably high risk of drug-induced liver injury.
@@
Against this backdrop, we argue that there is a correctable fundamental mismatch between the way LLMs work and the way chemists traditionally synthesize and study small molecules. Reducing tokenization granularity to the level of individual letters is counterproductive, as it complicates meaning extraction and increases the likelihood of generative AI hallucinate words that don’t exist. Analogous limitations are inherently linked to atom-based tokenization. And as a result, unfortunately, much of the generative AI research for scientific discovery within the computer science community does not extend to hypothesis verification in the physical world. To bridge this gap, in this paper, we propose a novel LLM by drawing inspiration from the scientific discovery process itself. We aim to develop a science-inspired large language model that follow three principles: (1) “Observe” - acquire, represent and integrate knowledge from multiple data modalities; (2) “Think” – think critically to generate hypotheses; and (3) “Propose” – verify hypotheses through the Physical World. We aim to teach computers to speak two complementary languages: one that represents molecular building blocks (i.e., subgraph structures) indicative of specific functions and compatible with automated modular assembly, and another that describes these functions in natural language (Figure 2). Unlike existing approaches that add such knowledge as a post hoc step, we develop a function- and synthesis-aware modular chemical language model (mCLM). Inspired by bilingual speakers who frequently “code-switch” (naturally and often switch between their two languages within the same message; Poplack, 2013), we propose a novel neural encoder that integrates molecular structure and natural language. mCLM incorporates both function- and synthesis-related knowledge into the small molecule tokenization process a priori. First, we tokenize small molecules at the level of building blocks (graph substructures) that are able to predict function and are, by design, flexible, multi-scale, fully compatible with automated modular small molecule synthesis. We use graph neural networks to encode each building block. We then extract natural language sentences from the literature that describe the molecule’s functions and chemical reactions and synthesis constraints of various molecules, and seamlessly insert the encoded building blocks alongside the corresponding entity names to form the training data, as illustrated in Figure 2. By reasoning on such functional building blocks, mCLM guarantees to generate efficiently synthesizable molecules thanks to recent progress in block-based chemistry, while also improving the functions of molecules in a principled manner. During model inference, mCLM enables these functional modules to be predictably and automatically assembled into molecules with desired functions.
@@ 2 The Modular Chemical Language Model
In natural language modeling, tokenization identifies common substrings (such as words and subwords), which carry richer semantic information than sequences of characters. Similarly in nature, most small molecules are composed primarily of connected building blocks  (Lehmann et al., 2018; Trobe and Burke, 2018). There is likewise a high degree of inherent modularity in many medicines and materials  (Ertl and Schuhmann, 2019; Arkan et al., 2020; Andrews et al., 1984; Vitaku et al., 2014).To leverage the phenomena in chemistry and borrow the spirit from natural language modeling, we propose mCLM, a multimodal model that jointly encodes and understands natural language and molecules based on synthesis-friendly building blocks instead of atoms. In Section 2.1, we introduce the concept of molecular building blocks as a chemical “vocabulary” and describe the tokenization process to obtain the library of building blocks. Then in Section 2.2 we describe the mCLM architecture and its training. Finally, we introduce the reasoning mechanism of mCLM which refines molecule design over multiple iterations in Section 2.3.
@@ 2.1 A Function-Infused and Synthesis-Friendly Vocabulary
During dataset construction, we combine two tokenization strategies to balance coverage and synthetic feasibility: a synthesis-guaranteed tokenizer and a rule-based tokenizer (Figure 3). The synthesis-guaranteed tokenizer disconnects the molecule only at bonds that can be formed by a predetermined small set of reactions that can be performed in an automated manner. Specifically, it only permits bond disconnections that correspond to reactions compatible with state-of-the-art automated synthesis platforms. These three bonds are: amide coupling, Suzuki-Miyaura coupling, and Buchwald-Hartwig coupling (see Figure 3 tokenization rules; Tyrikos-Ergas et al., 2025). The disconnection rules are adapted from well-established principles in block-based and automated chemistry (Blair et al., 2022b; Li et al., 2015b). Qualitative examples of tokenization are included in the appendix F, and the full rules and tokenizer code will be released with documentation. Importantly, the tokenizer is modular: as new automated reactions are validated in practice, their corresponding disconnection rules can be readily incorporated, expanding the makeable chemical space without requiring fundamental architectural changes. Unlike traditional retrosynthesis tools that prioritize maximum coverage or human-designed heuristics, this tokenizer is guided by the operational constraints of block chemistry. It also ensures that resulting blocks are free from functional group conflicts that would interfere with downstream synthesis. This conservative approach leads to a set of molecular tokens that are guaranteed to work under defined synthesis protocols, enabling seamless transition from model-generated output to real-world synthesis. When the synthesis-guaranteed tokenizer cannot fully cover a molecule—due to chemical incompatibilities or reaction constraints—we fall back on a rule-based tokenizer to ensure coverage of a larger variety of molecules. The rule-based tokenizer also breaks molecules along the same automated machine-friendly bonds as the above tokenizer. However, no other rules are applied beyond specifying a minimum size for blocks. This rule-based tokenizer is used only during training to support learning on diverse molecular structures while maintaining consistency with synthesis logic. We note that the vocabulary growth is analogous to the shift from character- level to subword-level tokenization in natural language, where larger but semantically meaningful tokens significantly improved performance.
@@ 2.2 Chemical-Language Modeling
We train mCLM on top of open-source pretrained large language models (we adopt Qwen2.5 Yang et al. (2024) as the starting point), allowing us to leverage their natural language understanding and reasoning capabilities without incurring the computational cost of training from scratch. As the training objective, we adopt a unified categorical cross-entropy (CCE) loss applied to both natural language and molecular tokens:
@@
Specifically, the cross-entropy loss is computed between the ground truth distribution $P(\mathbf{x})$ and the model distribution $P_{\theta}(\mathbf{x})$ over the combined vocabulary of natural language and molecular building blocks. The model generates logits for the next token $v$ by computing the dot product between the contextual representation $\mathbf{c}_{i}$ with token embeddings. $\mathbf{c}_{i}$ is produced by the Transformer given the previous tokens $\mathbf{x}_{1\cdots i-1}=[X_{1},\cdots,X_{i-1}])$. For natural language tokens, the embedding $\mathbf{e}_{v}$ is directly taken from the pretrained natural language model. The embedding for molecular building blocks is computed by passing the building block’s graph through a GNN, followed by a linear adapter function $f_{\psi}$ to project it into the same embedding space. This formulation enables joint training over both modalities using a single loss function. We train on approximately 1 million examples consisting of the 1,000 most frequent molecular building blocks, sampled from a new dataset we constructed. This dataset contains over 36 million molecular instructions, 1,000 tasks, and 8 million molecules. More details on the training procedure are described in Appendix E. Notably, the architecture of mCLM is compatible with larger vocabularies, and we anticipate future scaling studies as computational resources permit.
@@ 3.2 Improving FDA-Approved Drugs with Out-of-Vocabulary Blocks
In practice, real-world drug discovery is often driven by the optimization of known molecules, as existing approved drugs offer a more direct and promising path to safe and effective new drugs. Therefore, we evaluate mCLM’s drug proposal capability in improving FDA-approved drugs. Specifically, we apply the mCLM to improve the 6 properties for all FDA-approved drugs consisting of synthesis-guaranteed blocks. This amounts to 122 molecular structures and 153 unseen blocks. We confine the output vocabulary of mCLM to 582 synthesis-guaranteed blocks. Most of these drugs (120/122) contain blocks that were not present in the 1,000-block training vocabulary, presenting an opportunity to examine how the mCLM works on out-of-distribution molecules. Results in Table 3.2 show that improvements are achieved for all properties [^2] We also test a different distribution of 430 non-synthesis-guaranteed FDA drugs in Appendix A.2.. This shows that mCLM is not limited to a fixed vocabulary but can generalize beyond its training coverage. We also compare our approach against a wide range of strong text-based molecule editing baselines, including MoleculeSTM (Liu et al., 2022), FineMolTex (Li et al., 2025b), GPT-4o, GPT-5, Gemini-2.5-Flash (Gemini-2.5-F), LDMol (Chang and Ye, 2024), and Claude 3.5 Haiku (Claude-3.5-H). Despite the fact that all of the baseline models are allowed to generate without synthesis guarantees, their average improvements fall behind mCLM.
@@
The key to expediting the drug creation process is to discover potent molecular candidates that are simultaneously synthesis-friendly. While mCLM shows strong property editing results, its key benefit lies in its synthesis-friendly nature. We assessed the synthesizability of generated molecules by computing synthetic accessibility (SA scores) (Ertl and Schuffenhauer, 2009) as a quick heuristic (see Table 3). Then, as a more rigorous assessment, we consider Allchemy , which is the state-of-the-art retrosynthesis software (Wołos et al., 2022; Strieth-Kalthoff et al., 2024c). Allchemy is computationally expensive, but it evaluates synthesizability to the best of publicly available human chemical knowledge. For example, it finds synthetic routes for 98.1% of the FDA-approved molecules. We select the top 3 models by SA score (mCLM, MolSTM, and FineMolTex) and randomly sample 200 generated molecules from each to be assessed by Allchemy on a supercomputing cluster.
Table 2: Synthetic accessibility (SA) (Ertl et al., 2009), validity, and retrosynthetic results across baselines. Synthesizability indicates a retrosynthetic route was found. Makeability =Valid $\times$ Synth.
Model SA Validity (%) Synthesizability (%) Makeability (%)
@@
As shown in Table 2, molecules proposed by mCLM are 100% valid (syntactically correct) and 98.2% synthesizable, which is superior even to the FDA drugs. In contrast, MoleculeSTM outputs are valid only 93.9% of the time, and among those, only 90.3% are predicted to be synthesizable using exhaustive retrosynthesis search. Out of MoleculeSTM-generated molecules, only 85.4% can be made ($93.8\times 0.91=85.4$).

### v3

# arXiv:2505.12565v3[cs.AI] 01 Mar 2026

@@ Abstract
Despite their ability to understand chemical knowledge, large language models (LLMs) remain limited in their capacity to propose novel molecules with desired functions (e.g., drug-like properties). In addition, the molecules that LLMs propose can often be challenging to make, and are almost never compatible with automated synthesis approaches. To better enable the discovery of functional small molecules, LLMs need to learn a new molecular language that is more effective in predicting properties and inherently synced with automated synthesis technology. Current molecule LLMs are limited by representing molecules based on atoms. In this paper, we argue that just like tokenizing texts into meaning-bearing (sub-)word tokens instead of characters, molecules should be tokenized at the level of functional building blocks, i.e., parts of molecules that bring unique functions and serve as effective building blocks for real-world automated laboratory synthesis. This motivates us to propose mCLM, a modular Chemical-Language Model that comprises a bilingual language model that understands both natural language descriptions of functions and molecular blocks. mCLM front-loads synthesizability considerations while improving the predicted functions of molecules in a principled manner. Experiments on 430 FDA-approved drugs showed that mCLM is capable of significantly improving chemical functions critical to determining drug potentials. mCLM, with only 3B parameters, also achieves improvements in synthetic accessibility relative to 7 other leading generative AI methods including GPT-5. When tested on 122 out-of-distribution medicines using only building blocks/tokens that are compatible with automated modular synthesis, mCLM outperforms all baselines in property scores and synthetic accessibility. mCLM can also reason on multiple functions and iteratively self-improve to rescue drug candidates that failed late in clinical trials (“fallen angels”).
[Uncaptioned image] Data and Model Code
@@ 1 Introduction
Therefore, our goal is not to make all drug-like molecules; rather, our goal is to figure out how to make the right ones, better, faster, stronger. Unlike machines, human scientists are inherently “multilingual,” seamlessly navigating diverse modalities—from natural language and scientific figures in literature to complex scientific data such as molecular structures in knowledge bases. In contrast, most prior work on molecule discovery trains large language models on a single modality. Moreover, Human scientists “think before they talk,” grounding their reasoning in deeply reflective and deliberate reflection and critical evaluation to generate new hypotheses. Current models lack this critical thinking capacity, limiting their ability to contribute meaningfully to discovery. In particular, the human body is a highly complex, interconnected system, and drug discovery is essentially a multi-objective optimization problem. It involves balancing factors such as drug absorption, first-pass metabolism, bioavailability, distribution, protein binding, and clearance. However, improving one property often comes at the expense of another. Many promising drug candidates have failed in the final stages of FDA approval due to an unacceptably high risk of drug-induced liver injury.
@@
Against this backdrop, we argue that there is a correctable fundamental mismatch between the way LLMs work and the way chemists traditionally synthesize and study small molecules. Reducing tokenization granularity to the level of individual letters is counterproductive, as it complicates meaning extraction and increases the likelihood of generative AI hallucinating words that don’t exist. Analogous limitations are inherently linked to atom-based tokenization. And as a result, unfortunately, much of the generative AI research for scientific discovery within the computer science community does not extend to hypothesis verification in the physical world. To bridge this gap, in this paper, we propose a novel LLM by drawing inspiration from the scientific discovery process itself. We aim to develop a science-inspired large language model that follow three principles: (1) “Observe” - acquire, represent and integrate knowledge from multiple data modalities; (2) “Think” – think critically to generate hypotheses; and (3) “Propose” – verify hypotheses through the Physical World. We aim to teach computers to speak two complementary languages: one that represents molecular building blocks (i.e., subgraph structures) indicative of specific functions and compatible with automated modular assembly, and another that describes these functions in natural language (Figure 2). Unlike existing approaches that add such knowledge as a post hoc step, we develop a function- and synthesis-aware modular chemical language model (mCLM). Inspired by bilingual speakers who frequently “code-switch” (naturally and often switch between their two languages within the same message; Poplack, 2013), we propose a novel neural encoder that integrates molecular structure and natural language. mCLM incorporates both function- and synthesis-related knowledge into the small molecule tokenization process a priori. First, we tokenize small molecules at the level of building blocks (graph substructures) that are able to predict function and are, by design, flexible, multi-scale, fully compatible with automated modular small molecule synthesis. We use graph neural networks to encode each building block. We then extract natural language sentences from the literature that describe the molecule’s functions and chemical reactions and synthesis constraints of various molecules, and seamlessly insert the encoded building blocks alongside the corresponding entity names to form the training data, as illustrated in Figure 2. By reasoning on such functional building blocks, mCLM guarantees to generate efficiently synthesizable molecules thanks to recent progress in block-based chemistry, while also improving the functions of molecules in a principled manner. During model inference, mCLM enables these functional modules to be predictably and automatically assembled into molecules with desired functions.
@@ 2 The Modular Chemical Language Model
In natural language modeling, tokenization identifies common substrings (such as words and subwords), which carry richer semantic information than sequences of characters. Similarly in nature, most small molecules are composed primarily of connected building blocks  (Lehmann et al., 2018; Trobe and Burke, 2018). There is likewise a high degree of inherent modularity in many medicines and materials  (Ertl and Schuhmann, 2019; Arkan et al., 2020; Andrews et al., 1984; Vitaku et al., 2014). To leverage this phenomenon in chemistry and borrow the spirit from natural language modeling, we propose mCLM, a multimodal model that jointly encodes and understands natural language and molecules based on synthesis-friendly building blocks instead of atoms. In Section 2.1, we introduce the concept of molecular building blocks as a chemical “vocabulary” and describe the tokenization process to obtain the library of building blocks. Then in Section 2.2 we describe the mCLM architecture and its training. Finally, we introduce the reasoning mechanism of mCLM which refines molecule design over multiple iterations in Section 2.3.
@@ 2.1 A Function-Infused and Synthesis-Friendly Vocabulary
During dataset construction, we combine two tokenization strategies to balance coverage and synthetic feasibility: a synthesis-guaranteed tokenizer and a rule-based tokenizer (Figure 3). The synthesis-guaranteed tokenizer disconnects the molecule only at bonds that can be formed by a predetermined small set of reactions that can be performed in an automated manner. Specifically, it only permits bond disconnections that correspond to reactions compatible with state-of-the-art automated synthesis platforms. These three bonds are: amide coupling, Suzuki-Miyaura coupling, and Buchwald-Hartwig coupling (see Figure 3 tokenization rules; Tyrikos-Ergas et al., 2025). The disconnection rules are adapted from well-established principles in block-based and automated chemistry (Blair et al., 2022b; Li et al., 2015b). Qualitative examples of tokenization are included in the appendix F, and the full rules and tokenizer code will be released with documentation. Importantly, the tokenizer is modular: as new automated reactions are validated in practice, their corresponding disconnection rules can be readily incorporated, expanding the makeable chemical space without requiring fundamental architectural changes. Unlike traditional retrosynthesis tools that prioritize maximum coverage or human-designed heuristics, this tokenizer is guided by the operational constraints of block chemistry. It also ensures that resulting blocks are free from functional group conflicts that would interfere with downstream synthesis. This conservative approach leads to a set of molecular tokens that are guaranteed to work under defined synthesis protocols, enabling seamless transition from model-generated output to real-world synthesis. When the synthesis-guaranteed tokenizer cannot fully cover a molecule—due to chemical incompatibilities or reaction constraints—we fall back on a rule-based tokenizer to ensure coverage of a larger variety of molecules. The rule-based tokenizer also breaks molecules along the same automated machine-friendly bonds as the above tokenizer. However, no other rules are applied beyond specifying a minimum size for blocks. This rule-based tokenizer is used only during training to support learning on diverse molecular structures while maintaining consistency with synthesis logic. We note that the vocabulary growth is analogous to the shift from character- level to subword-level tokenization in natural language, where larger but semantically meaningful tokens significantly improved performance. These reusable, synthesis-friendly building blocks are derived from real data distributions. We collected approximately 8 million molecules from a wide range of public sources (Appendix G), of which approximately 6.8 million were tokenizable using our full tokenization pipeline (see Figure 3). This process resulted in a vocabulary of  800,000 unique building blocks with  200,000 synthesis-guaranteed (see Table 6). An example of how the block sequence can reconstruct the molecule structure is also provided in Appendix F.
@@ 2.2 Chemical-Language Modeling
We train mCLM on top of open-source pretrained large language models (we adopt Qwen2.5 Yang et al. (2024) as the starting point), allowing us to leverage their natural language understanding and reasoning capabilities without incurring the computational cost of training from scratch. Furthermore, pretrained natural-language LLMs have been exposed to a broad range of scientific information, making them a strong foundation for our curriculum. They help the model acquire general scientific knowledge before progressing to domain-specific training on molecular structures and drug-related information. As the training objective, we adopt a unified categorical cross-entropy (CCE) loss applied to both natural language and molecular tokens:
@@
Specifically, the cross-entropy loss is computed between the ground truth distribution $P(\mathbf{x})$ and the model distribution $P_{\theta}(\mathbf{x})$ over the combined vocabulary of natural language and molecular building blocks. The model generates logits for the next token $v$ by computing the dot product between the contextual representation $\mathbf{c}_{i}$ with token embeddings. $\mathbf{c}_{i}$ is produced by the Transformer given the previous tokens $\mathbf{x}_{1\cdots i-1}=[X_{1},\cdots,X_{i-1}])$. For natural language tokens, the embedding $\mathbf{e}_{v}$ is directly taken from the pretrained natural language model. The embedding for molecular building blocks is computed by passing the building block’s graph through a GNN, followed by a linear adapter function $f_{\psi}$ to project it into the same embedding space. This formulation enables joint training over both modalities using a single loss function. We train on approximately 1 million examples consisting of the 1,000 most frequent molecular building blocks, sampled from a new dataset we constructed. This dataset contains over 36 million molecular instructions, 1,000 tasks, and 8 million molecules. More details on the training procedure are described in Appendix E. Notably, the architecture of mCLM is compatible with larger vocabularies, and we anticipate future scaling studies as computational resources permit. mCLM is able to generalize to molecules composed of unseen building blocks. During training, the model learns to associate chemical structure with function via instruction-based pretraining. During inference, new blocks can be directly incorporated without retraining.
@@ 3.2 Improving FDA-Approved Drugs with Out-of-Vocabulary Blocks
In practice, real-world drug discovery is often driven by the optimization of known molecules, as existing approved drugs offer a more direct and promising path to safe and effective new drugs. Therefore, we evaluate mCLM’s drug proposal capability in improving FDA-approved drugs. Specifically, we apply the mCLM to improve the 6 properties for all FDA-approved drugs consisting of synthesis-guaranteed blocks. This amounts to 122 molecular structures and 153 unseen blocks. We confine the output vocabulary of mCLM to 582 synthesis-guaranteed blocks. Most of these drugs (120/122) contain blocks that were not present in the 1,000-block training vocabulary, presenting an opportunity to examine how the mCLM works on out-of-distribution molecules. Results in Table 3.2 show that improvements are achieved for all properties [^1] We also test a different distribution of 430 non-synthesis-guaranteed FDA drugs in Appendix A.2.. This shows that mCLM is not limited to a fixed vocabulary but can generalize beyond its training coverage. We also compare our approach against a wide range of strong text-based molecule editing baselines, including MoleculeSTM (Liu et al., 2022), FineMolTex (Li et al., 2025b), GPT-4o, GPT-5, Gemini-2.5-Flash (Gemini-2.5-F), LDMol (Chang and Ye, 2024), and Claude 3.5 Haiku (Claude-3.5-H). We conducted additional comparisons with two relevant baselines: DGAE (Boget et al., 2024) (a discrete graph autoencoder using vector quantization) and HierVAE (Jin et al., 2020) (a hierarchical VAE for molecular graph generation). The detailed results and a technical comparison with vector-quantized methods are enclosed in Appendix H. Despite the fact that all of the baseline models are allowed to generate without synthesis guarantees, their average improvements fall behind mCLM.
@@
The key to expediting the drug creation process is to discover potent molecular candidates that are simultaneously synthesis-friendly. While mCLM shows strong property editing results, its key benefit lies in its synthesis-friendly nature. We assessed the synthesizability of generated molecules by computing synthetic accessibility (SA scores) (Ertl and Schuffenhauer, 2009) as a quick heuristic (see Table 3). Then, as a more rigorous assessment, we consider Allchemy , which is the state-of-the-art retrosynthesis software (Wołos et al., 2022; Strieth-Kalthoff et al., 2024c). Allchemy is computationally expensive, but it evaluates synthesizability to the best of publicly available human chemical knowledge. For example, it finds synthetic routes for 98.1% of the FDA-approved molecules. Moreover, to our knowledge it is the only retrosynthesis software for which many of the proposed routes have been reduced to practice in the lab with physical experimentation. These studies have been published in top tier journals (Mikulak-Klucznik et al., 2020). We select the top 3 natural language models by SA score (mCLM, MolSTM, and FineMolTex) and randomly sample 200 generated molecules from each to be assessed by Allchemy on a supercomputing cluster.
Table 2: Synthetic accessibility (SA) (Ertl et al., 2009), validity, and retrosynthetic results across baselines. Synthesizability is the percent of valid molecules where a retrosynthetic route was found. Makeability is the overall percent of generations which can be synthesized (Makeability =Valid $\times$ Synth.).
Model SA $(\downarrow)$ Validity (%) Synthesizability (%) Makeability (%)
@@
mCLM (No GNN) 2.97 45.21 94.66 42.8
mCLM (No Synth. Tokenizer) 3.09 100.0 73.65 73.65
As shown in Table 2, molecules proposed by mCLM are 100% valid (syntactically correct) and 98.2% synthesizable, which is superior even to the FDA drugs. In contrast, MoleculeSTM outputs are valid only 93.9% of the time, and among those, only 90.3% are predicted to be synthesizable using exhaustive retrosynthesis search. Out of MoleculeSTM-generated molecules, only 84.8% can be made ($93.9\times 0.903=84.8$).
There are many new drug candidates that almost reach FDA approval but fall short for various reasons when tested in clinical trials. For example, Evobrutinib is a Bruton’s tyrosine kinase (BTK) inhibitor that went through clinical trial as a drug for relapsing Multiple Sclerosis. However, the FDA placed a partial clinical hold on Phase III trials in April 2023 after two patients showed signs of drug-induced liver injury (Montalban et al., 2024). TNG348 is a USP1 inhibitor designed for treating BRCA1/2-mutant and HRD cancers, but it failed in phase 1/2 clinical trials due to liver abnormalities (Inc, 2024; Simoneau et al., 2025). These “fallen angels” represent tremendous opportunities for impactful engagement of the AI/chemistry interface, because much is known about the strengths of each of these small molecules, and it is also known why they fell short. Fixing such fallen angels is a high leverage opportunity for the function-infused mCLM to contribute.
Figure 4: Examples of fallen angel property modification.
Figure 4 shows an application of the mCLM, without synthesis restrictions on the vocabulary, to these two molecules. For both, the initial step is to optimize DILI, the reason the drugs failed in clinical trials. Following that, the mCLM fixes other properties which were made worse in the previous attempt (PGP for Evobrutinib and BBBP for TNG248). For good measure, another property of each molecule is then improved. Notably, at each step, the mCLM only makes minor modifications to each drug of roughly 1 building block. While the mCLM shows promising results for repairing these drugs, it is worth noting that drug discovery is a many-objective optimization problem. While we are able to generate molecules with improved toxicity relative to Evobrutinib and TNG348, as well as other properties, yet other important properties may still have been compromised. Future work may want to investigate longer reasoning chains across a wider variety of properties. For comparison, we show a (less-successful) version of this experiment using MoleculeSTM (Appendix A.3).